OpenStudy (anonymous):

Solve the equation. 4(2m – 1) = 16 m = 52 m = 178 m = 103 m = 4

OpenStudy (anonymous):

m = 5/2, m = 17/8, m = 10/3, m = 4

OpenStudy (nikato):

first, ditribute like a(b-c)=ab-ac

OpenStudy (anonymous):

M=5/2 4(2m-1)=16 8m-4=16 8m=20 m=2.5 or 5/2